OpenStudy (anonymous):

If cos a= 12/13 and sin b= 4/5, then sin(a-b) equals .............. Anyone willing to help? I got -.0072 but I'm not sure if I'm correct, please help

OpenStudy (campbell_st):

well you need to draw the triangles and find the missing sides so you get the right ratio values  then its sin(a - b) = sin(a)cos(b) - cos(a)sin(b) = 5/13*3/5 - 12/13*4/5 = 15/65 - 48/65 = - 33/65 I'd leave it as an exact value..